OpenStudy (anonymous):

20g of a radiioactive substance reduces to 2.5g in 2079 day. What is the half life nd decay constant of the substance?

OpenStudy (rogue):

20 g ---> 10 g ---> 5 g ---> 2.5 g That's 3 half lives. So 1 half life is 2079/3 = 693 days. Now use the equation relating half life & the decay constant.\[k = \frac {\ln 2}{t_{1/2}} = \frac {\ln 2}{693} \approx 0.00100021238 = 1.00 \times 10^{-3}\]